# NeRC: Neural Ranging Correction through Differentiable Moving Horizon Location Estimation

### v1

@@ 5.1. Sequential Data Loader
Drawing inspiration from the techniques of random sampling and sequential partitioning used for reading long sequence data in natural language processing (Zhang et al., 2023), we design a sequential data loader tailored for handling raw GNSS measurements, as illustrated in Figure 8. Long sequences from the data files are first truncated using random offsets and subsequently divided into equal-length subsequences. These subsequences serve as the fundamental units, which are then randomly assembled into mini-batches. Finally, a sliding window moves along the batched subsequences to perform moving horizon estimation.
@@ 5.2. MLP Configuration
This picture shows the sequential data loader for training NeRC.
@@ 5.3. MHE Configuration
The differentiable MHE is empowered by Theseus, a DNLS solver built for PyTorch (Pineda et al., 2022). The Gauss-Newton optimizer is selected as the optimization kernel, with the maximum number of iterations and step size empirically tuned for optimal performance. Specifically, we set the step size to 0.5 and limit the number of iterations to 10. To balance the trade-off between training performance and efficiency, as indicated by 6(d), we eliminate the arrival cost and set the horizon size to 15. During inference, we employ the MHE with the arrival cost and set the horizon size to 5.
@@ 5.4. Building Euclidean Distance Cost Maps
The flowchart to construct EDF cost maps
The workflow of constructing an EDF cost map is illustrated by Figure 9. We start with collecting raw reference street data using Google My Maps, where walking, biking, and driving routes are freely available. Such data can be exported as a file of waypoints in the Keyhole Markup Language (KML) format. However, the raw street waypoints are too sparse to provide efficient supervision. Thus, we leverage the cubic spline to interpolate these raw route waypoints to obtain a dense and smooth reference route. Then, we grid the interpolated map on the longitude-latitude lattice with a tunable resolution. The grid cells occupied by the passable route are set to 1, while the blocked cells are set to 0. In this way, the gridded map turns into a binary image for efficient Euclidean distance transform (EDT) computation. Many mature solutions to computing the EDT of binary images are available (Paglieroni, 1992; Breu et al., 1995; Maurer et al., 2003; Felzenszwalb and Huttenlocher, 2012). We use the bwdist method provided by MATLAB to compute the Euclidean distance field (EDF) of the gridded map. To make the EDF cost map more differentiable, we apply to the map matrix a Gaussian lowpass filter of size 5 with a standard deviation of 1 (Yang et al., 2023b). Then, we compute the loss function (18) using PyTorch’s grid_sample method.
@@ 6.1. Datasets
This picture displays the environment where our MobileSG data was collected. The reference street data and corresponding EDF cost maps are also provided.
@@
The visualization of datasets. The training and testing data were collected along the same routes at approximately the same time intervals on different dates.
@@
GSDC 2022: Google enhanced the quality of the datasets in the second year of the competition, particularly in terms of ground truth calibration (Fu et al., 2022). However, the dataset only contains trajectories in open areas, where ranging measurement errors are trivial. Additionally, most traces in this dataset are annotated with only 2D location labels. We extract from it two subsets along two trajectories: GSDC22-1 and GSDC22-2.
@@
MobileSG: We collected a dataset in extremely challenging urban canyon areas using various Android devices, as illustrated in Figure 10. Ground truth locations were obtained only for the testing data, while the training data remained unlabeled. During testing data collection, the user’s movement was continuously recorded by a camera and later aligned with a 3D map in Google Earth Pro to determine the user’s locations (Liu et al., 2018b). This annotation method can achieve sub-meter accuracy (Sarlin et al., 2023), which is sufficient given that positioning errors in urban environments can reach up to several hundred meters (Hsu, 2018). As shown in Figure 10, the EDF cost map can be computed using publicly available street information corresponding to each training trajectory. The MobileSG dataset is used exclusively for evaluating the performance of EDF-based training.
@@ 6.2. Baseline Methods
We also compared NeRC with SOTA end-to-end learning approaches, including the black-box GnssFormer and the structured learning paradigms E2E-PrNet (Weng et al., 2024c) and pyrtklib (Hu et al., 2025; Hernández Olcina et al., 2024a). GnssFormer is our enhanced adaptation of the Set Transformer (Kanhere et al., 2022), incorporating the same input features used in NeRC and initialized with robust location estimates to guide its position corrections. Both E2E-PrNet and pyrtklib adopt the structure composed of an MLP followed by a differentiable WLS positioning engine. E2E-PrNet also leverages the Theseus library for differentiable optimization, while pyrtklib integrates a differentiable Python binding of the popular rtklib library. We have enhanced these three baseline approaches with our proposed 2D and EDF-based training losses, enabling them to handle partially labeled or unlabeled data like NeRC. All the baseline methods we implemented or trained will be publicly available at: https://github.com/AILocAR/NeRC.
Horizontal positioning results.
6.3. Results
Evaluation metrics: Most location-based mobile applications require only the device’s horizontal position. In the GSDC competitions, evaluation is conducted solely on the horizontal localization error, which is measured by the horizontal distance between the predicted and ground truth latitude/longitude coordinates. Accordingly, we evaluate only horizontal positioning performance in this work. The horizontal distance is calculated using Vincenty’s formulae (Vincenty, 1975), as implemented in the pymap3d library. 12(h) illustrates the boxplots of horizontal errors of all the methods previously mentioned. Additionally, in accordance with the GSDC competition requirements, we report the horizontal score for each method, indicated by the red dot in 12(h) and summarized in Table 2. The horizontal score is defined as the average of the 50th and 95th percentiles of the horizontal errors. A lower horizontal score indicates higher positioning accuracy.
Positioning performance: 12(h) and Table 2 indicate that NeRC outperforms all baseline methods across all datasets.
(1) Large-scale erroneous scenarios. GSDC21R represents a large-scale erroneous scenario. As illustrated in Figure 11a, the data were collected along 120 km of open highways in the Bay Area, yet exhibit positioning errors of approximately 15 meters, as reflected in the model-based method results in 12(a). Despite the open-sky environment, such inaccuracies may stem from system-level issues, particularly ground truth calibration errors (Fu et al., 2022). In this case, all data-driven methods perform well, with NeRC achieving horizontal accuracy comparable to FCNN-LSTM and PrNet.
(2) Large-scale open scenarios.
GSDC22 and GSDC23 were collected in large-scale open-sky scenes where model-based baselines already achieve high localization accuracy (within 5 meters), owing to strong signal quality. In such environments, ranging errors are trivial and often comparable to the inherent measurement noise. As a result, methods like E2E-PrNet and pyrtklib tend to overfit the noise, leading to degraded performance. In contrast, NeRC, supported by differentiable MHE, effectively suppresses noise and corrects residual ranging errors. GnssFormer, which directly regresses location corrections in a black-box fashion, is sensitive to its initialization quality—here, EKF estimates—around which its outputs fluctuate (Mohanty and Gao, 2023; Weng et al., 2024b). With fully labeled data, PrNet and FCNN-LSTM can be trained on filtered ranging errors (Weng et al., 2024b), which explains their performance approaching that of NeRC in 12(c). However, their reliance on complete 3D labeled data precludes their use on GSDC22, where only 2D location labels are available.
(3) Small-scale urban canyons. GSDC21U and MobileSG represent some of the most challenging small-scale urban localization scenarios. Despite severe multipath, NLOS effects, and high measurement noise, NeRC consistently outperforms all competing methods, achieving accuracy gains in both fully labeled and unlabeled settings. On the fully labeled GSDC21U dataset, NeRC surpasses the best model-based (MHE) and data-driven (PrNet) baselines by approximately 65% and 49%, respectively. On the unlabeled MobileSG dataset, NeRC improves over the leading model-based method (FGO) by 16–31% and outperforms data-driven approaches (E2E-PrNet and GnssFormer) by 12–36%. Therefore, NeRC reaches the SOTA level and maintains strong localization performance under these challenging conditions.
@@ 6.3. Results
[…]
@@
Impact of Training Methods: 13(a) compares the effects of different training methods on the positioning errors of NeRC. The MHE algorithm is used as the baseline. The same MHE is also integrated into NeRC as the downstream localization engine to demonstrate NeRC’s improvement. The evaluation is conducted on GSDC21U. As shown 13(a), incorporating the EDF cost map can help reduce positioning errors without using ground-truth location data. Training with either 2D or 3D labels yields comparable performance, both contributing to further improvements in NeRC’s localization accuracy.
Overhead of Training and Inference:
As illustrated in 13(b), we evaluate the computational time and GPU memory usage of end-to-end learning-based approaches during both training and inference. Compared to existing SOTA methods, our proposed NeRC generally introduces higher computational overhead, as it processes a horizon of measurements simultaneously. Among these methods, GnssFormer achieves the fastest training and inference speed, benefiting from PyTorch’s highly optimized Transformer implementation and its precomputed reference trajectories. In contrast, E2E-PrNet and NeRC exhibit longer running time due to repeatedly computing positions with a relatively less optimized nonlinear least squares solver (Pineda et al., 2022). Although pyrtklib leverages a standard MLP and a basic WLS algorithm, without involving any nonlinear solver, it incurs an even longer training time because it currently only supports sequential training (Hu et al., 2025). About memory usage, although NeRC requires more resources during training, its memory consumption during inference is comparable to that of other methods. Its real-time performance will be further analyzed in section 7.
These pictures show the impact of training loss and the computational cost of different end-to-end methods.
The edge-based NeRC system for mobile localization is a typical client-server architecture.
@@ 7. Real-Time Field Test
One potential real-world application of NeRC is to provide ranging correction services over a network when users enter GNSS-challenged environments (Liu et al., 2018b; Hernández Olcina et al., 2024b). To this end, we design an edge-based mobile localization system powered by NeRC and evaluate its real-time performance in urban canyon areas characterized by dense high-rise buildings.
@@
System Implementation: On the client side, we built the Android app in Java using Android Studio based on Google’s open-source software GnssLogger (Fu et al., 2020). Raw GNSS measurements are collected with a frequency of 1 Hz via the Android Location API. Then, the collected data are packaged into JSON files for communicating with the edge server through the HTTP protocol (Scargill et al., 2023). The app runs on an Android 13 phone powered by the Snapdragon 8+ Gen 1 processor with 12 GB RAM. Only the GPS L1 signals are utilized in our experiment.
These figures depict the physical real-time field test, which is set by an edge server, a WAP, and a user mobile application. The experiment is conducted in a challenging scenario where multipath/NLOS conditions are severe.
The edge server is implemented using Uvicorn, an asynchronous web framework in Python, to listen and respond to requests from the Android client (Duan et al., 2024). The received data are cached in a moving horizon, the size of which is tuned to 5. If the interval between the timestamps of the latest two messages is greater than 10 seconds, the data horizon will be considered discontinuous and reset to only hold the current message. The NeRC framework is implemented and trained in the same way as in the previous section, but is just running in inference mode (Hu et al., 2024). The server runs Ubuntu 18.04.6 LTS, powered by an Intel Core i7-10750H 2.6 GHz CPU with 16 GB RAM and an NVIDIA GeForce RTX 2060 GPU.
15(c) illustrates the physical setup of the real-time test. The experiment was conducted in a densely built urban area, where a road was surrounded by high-rise buildings, as shown in 15(a). A user holding an Android phone walked along the road and localized himself using the NeRC system. As shown in 15(b), a 2.4 GHz wireless access point was installed on a second-floor balcony to provide the widest possible coverage and to connect the server and the user within the same WLAN. Our Android app, displayed in 15(c), enables communication with the server once the user inputs the server’s IP address.
@@
(1) Localization accuracy. The reference trajectory is scheduled ahead of time, as shown in 16(a). The user walked along the reference route, and the corresponding positioning result is presented in 16(b). The red trajectory in 16(b) represents Google’s baseline solution provided by the open-source GnssLogger software (Fu et al., 2020). Note that this is not the positioning result from the official Google Maps app, which utilizes in-phone GNSS chipset locations computed via the manufacturer’s proprietary algorithms. The baseline trajectory is visibly noisy and biased due to the severe multipath/NLOS propagation effects caused by the surrounding high-rise buildings. In contrast, NeRC’s solution, depicted in blue, closely follows the reference trajectory, demonstrating higher robustness and accuracy.
These figures depict the physical real-time field test result. The experiment is conducted in a deep urban area. We walked along a pre-scheduled path and recorded the location trajectory from the mobile phone.
@@
(3) GPU overhead on the edge server. The edge server running NeRC only performs in the inference mode, and the deployed neural model is trained offline in advance. We trigger “nvidia-smi” programmatically to log the usage of the server’s GPU and its memory per second. The average usage of GPU memory across 123 samples is 1.4 GB, while the free GPU memory is up to 4.5 GB. Moreover, the average utilization of the GPU is just 4.8%. Thus, our NeRC model is lightweight enough to run on common consumer-level GPUs, such as the NVIDIA RTX 2060 we used.
(4) Computational latency on the edge server. We logged the inference time of NeRC on the edge server to evaluate its computational efficiency. On average, computing a location from a single epoch of measurements takes 100 ms, with a median latency of 98 ms. The inference time is closely influenced by the size of the neural network, the horizon length used in the MHE, and the performance of the GPU. Using a smaller neural network or horizon size on a more powerful GPU can reduce the latency.
(5) End-to-end latency: We define the end-to-end latency as the time interval required for the app to send raw measurements to the edge server and receive the localization results in return (Duan et al., 2024; Ding and Chen, 2025). Accordingly, we measured this latency on the client side. The end-to-end latency includes both the server-side computation time and the communication delay over the WLAN. In our experiment, the mean end-to-end latency was 246 ms, and the median was 225 ms. Given that our system operates at 1 Hz (i.e., Android collects raw GNSS measurements once per second), this level of latency has a negligible impact on the system’s real-time performance.
The CPU and memory usage are profiled by “View Live Telemetry” in Android Studio.

### v2

@@ 5.1. Sequential Data Loader
Drawing inspiration from the techniques for loading long sequences in natural language processing (Zhang et al., 2023), we design a sequential data loader tailored for handling raw GNSS measurements, as illustrated in Figure 8. Long sequences from data files are first truncated using random offsets and subsequently divided into equal-length subsequences. These subsequences serve as fundamental units, which are then randomly assembled into mini-batches. Finally, a sliding window moves along the batched subsequences to perform MHE.
@@ 5.3. MHE Configuration
The differentiable MHE is empowered by Theseus, a DNLS solver built for PyTorch (Pineda et al., 2022). The Gauss-Newton optimizer is selected as the optimization kernel, with the maximum number of iterations and step size empirically tuned for optimal performance. Specifically, we set the step size to 0.5 and limit the number of iterations to 10. To balance the trade-off between training performance and efficiency, as indicated by Figure 6, we eliminate the arrival cost and set the horizon size to 15. During inference, we employ the MHE with the arrival cost and set the horizon size to 5.
@@ 5.4. Building Euclidean Distance Cost Maps
The workflow of constructing an EDF cost map is illustrated by Figure 9. We start with collecting raw reference street data using Google My Maps, where walking, biking, and driving routes are freely available. Such data can be exported as a file of waypoints in the Keyhole Markup Language (KML) format. However, the raw street waypoints are too sparse to provide efficient supervision. Thus, we leverage the cubic spline to interpolate these raw route waypoints to obtain a dense and smooth reference route. Then, we grid the interpolated map on the longitude-latitude lattice with a tunable resolution. The grid cells occupied by the passable route are set to 1, while the blocked cells are set to 0. In this way, the gridded map turns into a binary image for efficient EDT. Many mature solutions to the EDT of binary images are available (Paglieroni, 1992; Breu et al., 1995; Maurer et al., 2003; Felzenszwalb and Huttenlocher, 2012). We use the bwdist method provided by MATLAB to compute the EDF of the gridded map. To make the EDF cost map more differentiable, we apply to the map matrix a Gaussian lowpass filter of size 5 with a standard deviation of 1 (Yang et al., 2023). Then, we compute the loss function (18) using PyTorch’s grid_sample method.
@@ 6.1. Datasets
GSDC 2022: Google enhanced the quality of their datasets in the second year of the competition, particularly in terms of ground truth calibration (Fu et al., 2022). However, the dataset only contains trajectories in open areas, where ranging measurement errors are trivial. Additionally, most traces in this dataset are annotated with only 2D location labels. We extract from it two subsets along two trajectories: GSDC22-1 and GSDC22-2.
@@
MobileSG: We collected a dataset in extremely challenging urban canyon areas using various Android devices, as illustrated in Figure 10. Ground truth locations were obtained only for the testing data, while the training data remained unlabeled. During testing data collection, the user’s movement was continuously recorded by a camera and later aligned with a 3D map in Google Earth Pro to determine the user’s locations (Liu et al., 2018b). This annotation method can achieve sub-meter accuracy (Sarlin et al., 2023), which is sufficient given that positioning errors in urban environments can reach up to several hundred meters (Hsu, 2018). As shown in Figure 10, the EDF cost maps are computed using publicly available street information corresponding to training trajectories. The MobileSG dataset is used exclusively for evaluating the performance of EDF-based training.
@@ 6.2. Baseline Methods
We also compared NeRC with SOTA end-to-end learning approaches, including the black-box GnssFormer and the structured learning paradigms E2E-PrNet (Weng et al., 2024c) and pyrtklib (Hu et al., 2025; Hernández Olcina et al., 2024a). GnssFormer is our enhanced adaptation of the Set Transformer (Kanhere et al., 2022), incorporating the same input features used in NeRC and initialized with robust location estimates to guide its position corrections. Both E2E-PrNet and pyrtklib adopt the structure composed of an MLP followed by a differentiable WLS positioning engine. E2E-PrNet also leverages the Theseus library for differentiable optimization, while pyrtklib integrates a differentiable Python binding of the popular rtklib library. We have enhanced these three baseline approaches with our proposed 2D and EDF-based training losses, enabling them to handle partially labeled or unlabeled data like NeRC.
In a dataset, the same labels and loss function are used to train all data-driven methods. The baseline methods we implemented or trained are released at https://github.com/AILocAR/NeRC.
@@
| 1 | 27.957 | 18.148 | 17.909 | 18.024 | ✗ | ✗ | 20.387 | 19.238 | 25.572 | 12.350 | ||
| 2 | 43.687 | 31.992 | 30.824 | 31.948 | ✗ | ✗ | 29.570 | 43.525 | 43.002 | 26.043 | ||
| 3 | 33.332 | 28.429 | 24.568 | 26.371 | ✗ | ✗ | 27.050 | 36.222 | 37.602 | 20.109 | ||
6.3. Results
Evaluation metrics: Most location-based mobile applications require only horizontal positioning. In the GSDC competitions, evaluation is based solely on horizontal localization errors, measured by horizontal distances between predicted and ground truth latitude/longitude coordinates. Accordingly, this work evaluates only horizontal positioning performance. The horizontal distance is calculated using Vincenty’s formulae (Vincenty, 1975), as implemented in the pymap3d library. Figure 12 illustrates boxplots of horizontal errors of all the methods previously mentioned. Additionally, in accordance with the GSDC competition requirements, we report horizontal scores for all methods, indicated by red dots in Figure 12 and summarized in Table 2. The horizontal score is defined as the average of the 50th and 95th percentiles of horizontal errors. A lower horizontal score indicates higher positioning accuracy.
Positioning performance: Figure 12 and Table 2 indicate that NeRC outperforms all baseline methods across all datasets.
(1) Large-scale erroneous scenarios. GSDC21R represents a large-scale erroneous scenario. As illustrated in Figure 11a, the data were collected along 120 km of highways in the Bay Area, yet exhibit positioning errors of approximately 15 meters, as reflected in the model-based method results in 12(a). Despite the open-sky environment, such inaccuracies may stem from system-level issues, particularly ground truth calibration errors (Fu et al., 2022). In this case, all data-driven methods perform well, with NeRC achieving horizontal accuracy comparable to FCNN-LSTM and PrNet.
(2) Large-scale open-sky scenarios.
GSDC22 and GSDC23 were collected in large-scale open-sky scenes where model-based baselines already achieve high localization accuracy, owing to strong signal quality. In such environments, ranging errors are trivial and often comparable to inherent measurement noise. As a result, methods like E2E-PrNet and pyrtklib tend to overfit to noise, leading to degraded performance. In contrast, NeRC, supported by differentiable MHE, effectively suppresses noise and corrects residual ranging errors. GnssFormer, learning location corrections in a black-box fashion, is sensitive to its initialization quality—here, EKF estimates—around which its outputs fluctuate (Mohanty and Gao, 2023; Weng et al., 2024b). PrNet and FCNN-LSTM are trained on filtered ranging error labels (Weng et al., 2024b), explaining their performance approaching NeRC in 12(c). However, their reliance on fully labeled 3D data precludes their use on GSDC22, where only 2D location labels are available.
(3) Small-scale urban canyons. GSDC21U and MobileSG represent some of the most challenging small-scale urban localization scenarios. Despite severe multipath, NLOS effects, and high measurement noise, NeRC consistently outperforms all competing methods, achieving accuracy gains in both fully labeled and unlabeled settings. On the fully labeled GSDC21U dataset, NeRC surpasses the best model-based (MHE) and data-driven (PrNet) baselines by approximately 65% and 49%, respectively. On the unlabeled MobileSG dataset, NeRC improves over the leading model-based method (FGO) by 16–31% and outperforms other end-to-end approaches (E2E-PrNet and GnssFormer) by 12–36%. Therefore, NeRC reaches the SOTA level and maintains strong localization performance under these challenging conditions.
@@ 6.3. Results
Impact of training label quality: Since NeRC is trained in a supervised manner, its performance is directly influenced by the accuracy of training labels. 13(a) compares the positioning errors of NeRC under different training strategies. The MHE algorithm serves as the baseline, and the same MHE is also integrated into NeRC as the downstream localization engine to demonstrate its enhancement. The evaluation is conducted on GSDC21U. Three EDF cost maps of varying qualities are provided for training with unlabeled data, reflecting different levels of prior knowledge about the routes used for data collection. As shown 13(a), incorporating the EDF cost map can help reduce positioning errors without using location ground truth. Moreover, higher-quality EDF cost maps lead to smaller errors. Training with either 2D or 3D labels yields comparable performance, both contributing to further improvements in NeRC’s localization accuracy.
Ranging error regression: 13(b) displays ranging errors at four epochs regressed by the upstream MLP, which was trained using 3D/2D ground truth locations and EDF cost maps. The reference ranging errors are calculated based on 3D location labels, as described in (Weng et al., 2024b). As illustrated in 13(b), at each epoch, the predicted ranging errors of visible satellites exhibit patterns similar to the reference, regardless of the training loss used. This indicates that the upstream MLP has successfully learned the underlying physical error patterns. For each kind of training loss, however, an offset nearly uniform across all satellites is observed between predictions and the reference, arising from missing label dimensions such as heights and device clock offsets (Weng et al., 2024c). Nevertheless, the shared bias among all satellites only influences timing estimation and does not affect positioning accuracy (Kaplan and Hegarty, 2017).
Generalization ability: Figure 14 illustrates the robustness of NeRC in generalizing to unseen paths during inference. In this experiment, inference data were collected along ten alleyways that were not included in the training routes. Results show that NeRC can still enhance GNSS positioning accuracy in several unseen alleyways, even though the model was originally designed to learn within a specific route. However, on certain paths, such as Path 6 and Path 7, NeRC yields degraded positioning performance compared to the traditional model-based MHE. These paths are fully surrounded by unseen tall buildings, resulting in a distribution of ranging errors that differs significantly from the training data.
@@ 7. Real-Time Field Test
One potential real-world application of NeRC is to provide ranging correction services over a network when users enter GNSS-challenged environments (Liu et al., 2018b; Hernández Olcina et al., 2024b). To this end, we design an edge-based mobile localization system powered by NeRC and evaluate its real-time performance in urban canyon areas.
@@
System Implementation: On the client side, we built the Android app in Java using Android Studio based on Google’s open-source software GnssLogger (Fu et al., 2020). Raw GNSS measurements are collected with a frequency of 1 Hz via the Android Location API. Then, the collected data are packaged into JSON files for communicating with the edge server through the HTTP protocol (Scargill et al., 2023). The app runs on an Android 13 phone powered by the Snapdragon 8+ Gen 1 processor with 12 GB RAM. Only the GPS L1 signals are utilized in our experiment. The edge server is implemented using Uvicorn, an asynchronous Python web framework, to listen for and respond to requests from the Android client (Duan et al., 2024). The size of the horizon caching received data is set to 5. If the interval between the timestamps of the latest two messages exceeds 10 seconds, the data horizon will be considered discontinuous and reset to hold only the current message. NeRC is trained using EDF cost maps and just runs in inference mode (Hu et al., 2024). The server runs Ubuntu 18.04.6 LTS, powered by an Intel Core i7-10750H 2.6 GHz CPU with 16 GB RAM and an NVIDIA GeForce RTX 2060 GPU.
Figure 16 illustrates the physical setup of the real-time test. The experiment was conducted in a densely built urban area, where a road was surrounded by high-rise buildings, as shown in 16(a). A user holding an Android phone walked along the road and localized himself using the NeRC system. As shown in 16(b), a 2.4 GHz wireless access point was installed on a second-floor balcony to provide the widest possible coverage and to connect the server and the user within the same WLAN. Our Android app, displayed in 16(c), enables communication with the server once the user inputs the server’s IP address.
@@
(1) Localization accuracy. The reference trajectory is scheduled ahead of time, as shown in 17(a). The user walked along the reference route, and the corresponding positioning result is presented in 17(b). The red trajectory in 17(b) represents Google’s baseline solution provided by the open-source GnssLogger software (Fu et al., 2020). Note that this is not the positioning result from the official Google Maps app, which utilizes in-phone GNSS chipset locations computed via the manufacturer’s proprietary algorithms. The baseline trajectory is visibly noisy and biased due to severe multipath/NLOS propagation effects from the surrounding high-rise buildings. In contrast, NeRC’s solution (depicted in blue) closely follows the reference trajectory, demonstrating higher accuracy and greater robustness to noise.
@@
(3) GPU overhead on the edge server. The edge server runs NeRC only in the inference mode. We trigger “nvidia-smi” programmatically to log the usage of the server’s GPU and its memory per second. The average usage of GPU memory across 123 samples is 1.4 GB, while the free GPU memory is up to 4.5 GB. Moreover, the average utilization of the GPU is just 4.8%. Thus, our NeRC model is lightweight enough to run on common consumer-level GPUs, such as the NVIDIA RTX 2060 we used.
(4) Computational latency on the edge server. We logged the inference time of NeRC on the edge server to evaluate its computational efficiency. On average, computing a location from a single epoch of measurements takes 100 ms, with a median latency of 98 ms. Inference time is closely influenced by the neural network size, the horizon length used in MHE, and the GPU’s performance. Using a smaller neural network or horizon size on a more powerful GPU can reduce the latency.
(5) End-to-end latency: We define the end-to-end latency as the time interval required for the app to send raw measurements to the edge server and receive localization results in return (Duan et al., 2024; Ding and Chen, 2025). Accordingly, we measured this latency on the client side. The end-to-end latency includes both the server-side computation time and the communication delay over the WLAN. In our experiment, the mean end-to-end latency was 246 ms, and the median was 225 ms. Given that our system operates at 1 Hz (i.e., Android collects raw GNSS measurements once per second), this level of latency has a negligible impact on the system’s real-time performance.